# Upper bounds on the highest phonon frequency and superconducting temperature from fundamental physical constants

### v1

@@ 1 Introduction
The governing role of fundamental physical constants has been extensively discussed in high-energy particle physics and astrophysics. Examples include the stability of nuclei, nuclear reactions, and the formation and evolution of stars – where heavy elements are produced and give rise to other observable structures. In these and other processes, fundamental constants give the Universe its observed properties and separate it from potential others Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Understanding the origin of fundamental constants and their fine-tuning is considered one of the grand challenges in science Barrow and Webb (2006); Weinberg (1983). Approaching this challenge starts with deepening our understanding of how fundamental constants affect phenomena spanning the length and energy scales between particle physics and astrophysics Trachenko (2023a); Brazhkin (2023); Trachenko (2023b).
@@
In condensed matter physics and in other areas that support a continuum description of matter, such as elasticity and hydrodynamics, system properties are driven by many-body collective effects. These effects are often considered as emergent and thus not reducible to individual particles whose properties are set by fundamental constants. It was therefore unexpected when it was discovered that fundamental constants govern the bounds, characteristic values and properties of several key many-body collective processes. Examples include viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), thermal expansion, melting Trachenko (2024) and velocity gradients that can be set up in cells using biochemical energy Trachenko (2023c). In these examples, properties are governed by either fundamental constants only or by a combination of these constants with other parameters such as temperature. The implications of these findings extend to both fundamental and applied science Trachenko (2023a).
@@
Here, we show that fundamental physical constants provide an upper bound to the frequency of phonons in condensed matter phases, setting a limit to the vibrational frequency of atoms. This bound involves the dimensionless electron-to-proton mass ratio $\beta$, the key constant involved in setting the finely-tuned “habitable zone” of our world in the ($\alpha$, $\beta$) space, where $\alpha$ is the fine-structure constant Barrow (2003). We find this bound to be in agreement with ab initio simulations of atomic hydrogen and hydride superconductors which, as hydrogen is the lightest nucleus (a single proton), host the highest possible phonon frequencies. We subsequently show that this upper bound in the maximal phonon frequency implies an upper bound to the phonon-mediated superconducting critical temperature $T_{\mathrm{c}}$ in condensed matter. In particular, we show that fundamental constants set the upper bound to $T_{c}$ on the order of $10^{2}-10^{3}$ K, consistent with calculations of $T_{c}$ from optimal Eliashberg functions. We also observe that in addition to electron-phonon coupling, the bound to $T_{c}$ in terms of fundamental constants applies to other coupling mechanisms such as electron-electron coupling. This implies that the current search for $T_{\mathrm{c}}$ at and above $300$ K is itself due to fundamental constants taking their observed values. We finally observe that fundamental constants affect the observability of other phenomena in condensed matter physics including phase transitions.
@@ 2 Upper bound on the vibrational frequency
and the Rydberg energy, $E_{\rm R}=\frac{e^{2}}{8\pi\epsilon_{0}a_{\rm B}}$ Ashcroft and Mermin (1976), sets the characteristic scale of $E$ to the order of several eV:
@@
The same ratio (3) follows by combining two known relations in metallic systems: $\frac{\hbar\omega_{\rm m}}{E}\approx\frac{c}{v_{\rm F}}$, where $v_{\rm F}$ is the Fermi velocity and $c$ is the speed of sound, and $\frac{c}{v_{\rm F}}\approx\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$, providing an order-of-magnitude estimation of $\frac{\hbar\omega_{\rm m}}{E}$ in other systems too Ashcroft and Mermin (1976).
@@
Equation  (5) offers a simple result showing that the upper bound to the atomic oscillation frequency is set by the Compton frequency times the square of $\alpha$ and the square-root of the electron-to-phonon mass ratio.
Equation  (4) sets $\omega_{\rm m}^{\rm u}$ in terms of fundamental physical constants. The constants $\hbar$, $m_{e}$, and $e$ determine the electromagnetic energy scale driving atomic oscillations. The proton mass $m_{p}$ is intrinsic to the atoms being driven and sets the scale (both in time, and space) of the oscillatory response. Interestingly, $\omega_{\rm m}^{\rm u}$ involves the electron-to-proton mass ratio $\beta=\frac{m_{e}}{m_{p}}$. This dimensionless constant is considered of fundamental importance: together with the fine-structure constant, it sets a finely-tuned habitable zone in ($\alpha$, $\beta$) space where it governs the stability of molecular structures, processes involved in igniting and fuelling stars (and thus the production of heavy elements), and other essential effects giving rise to our observable world Barrow (2003).
@@
Equations  (4)-(6) show that fundamental constants set $\omega_{\rm m}^{\rm u}$ on the order of $10^{3}$$10^{4}$ K, a result we will use to discuss phonon-mediated superconductivity below. The upper bound $\omega_{\rm m}^{\rm u}$ in (6) is of the same order of magnitude, but above, the experimental Debye frequency $\omega_{\rm D}$ of 2240 K in ambient pressure diamond Tohei et al. (2006), the hardest known material Brazhkin and Solozhenko (2019).
The upper bound (4)-(6) corresponds to solid hydrogen with strong metallic bonding. Although this phase only exists at megabar pressures Dias and Silvera (2017); Loubeyre et al. (2020); McMahon et al. (2012) and is thermodynamically and dynamically unstable at ambient pressure, it is interesting to check the validity of our bound and calculate $\omega_{\rm m}$ in atomic hydrogen. The importance of this calculation is also highlighted by the strong interest in the properties of atomic hydrogen at high pressure (see, e.g., Refs. Dias and Silvera (2017); Loubeyre et al. (2020); McMahon et al. (2012)).
Before calculating $\omega_{\rm m}$, we first evaluate the effect of pressure on the fundamental bound (4) because solid atomic hydrogen exists at high pressure only as mentioned above. We make two remarks in relation to the pressure effect. First, hydrogen is a unique element with no core electrons. As compared to heavier elements, this gives weaker repulsive contributions to the interatomic interaction and weaker pressure dependence of elastic moduli Brazhkin and Lyapin (2002). As a result, pressure has a weaker effect on $\omega_{\rm m}$ in hydrogen as compared to other systems. Second, the effect of pressure on $\omega_{\rm m}^{\rm u}$ can be approximately estimated by adding the term $PV$ to $E$ in Eq. (3), representing the work needed to overcome the elastic deformation Frenkel (1947) due to external pressure in order to surmount the energy of cohesion. Here, $V$ is the elementary volume on the order of $a^{3}$. According to Eqs. (3)-(4), this increases $\omega_{m}^{u}(P)$ as
@@
We calculate the phonon dispersion curves of atomic hydrogen in the $I4_{1}/amd$ structure Nagao et al. (1997); Pickard and Needs (2007), which is currently the best candidate structure for solid atomic metallic hydrogen. This structure is calculated to become thermodynamically stable in the pressure range 400400400400600600600600 GPa Azadi et al. (2014); McMinis et al. (2015); Monacelli et al. (2023), below which solid hydrogen is a molecular solid. We perform density functional theory (DFT) calculations using the castep package Clark et al. (2005). Details of DFT calculations are given in the Appendix.
@@
Figure 2: Maximal phonon frequency as a function of pressure for different systems. Blue bullet points show the calculated maximal phonon frequencies in atomic hydrogen at the corresponding pressures. Red diamonds show the maximal phonon frequencies in a range of hydrides at different pressures, which are, from left to right: MgIrH6 Doliu et al. (2024), MgVH6 Zheng et al. (2021), CaH6 Jeon et al. (2022), LaH10 (top) Durajski et al. (2020), H3S (bottom) Sano et al. (2016), and SrH10 Tanaka et al. (2017). The horizontal solid black line shows the fundamental bound $\omega_{\rm m}^{\rm u}$ (6). The horizontal dash-dotted line shows the Debye frequency of diamond at ambient pressure. The dashed line on top of the graph shows $\omega_{\rm m}^{\rm u}(P)$ in Eq. (7) with $a=2a_{\rm B}$.
@@
We observe in Fig. 2 that the calculated $\omega_{\rm m}$ is only slightly above $\omega_{\rm m}^{\rm u}$ at the high pressure of 400 GPa in atomic hydrogen. Assuming a pressure-independent effective $a$ in Eq. (7) for estimation purposes, we find that $\omega_{\rm m}^{\rm u}(P)$ lies close and above the hydrogen points if $a$ is taken to be about $2a_{\rm B}$ (see dashed black line in Fig. 2). This is consistent with the elementary volume being related to the Bohr radius and fundamental constants. Although $a=2a_{\rm B}$ in Eq. (7) introduces an ad hoc choice, the dashed line in Fig. 2 demonstrates that the pressure effect on $\omega_{\rm m}^{\rm u}$ can be accounted for with a sensible and physically justified choice of $a$. In this regard, we note more generally that evaluating bounds on many important condensed matter properties using fundamental constants only inevitably involves approximations Trachenko (2023a); Brazhkin (2023). This includes $\omega_{\rm m}^{\rm u}$ in Eqs. (4)-(6) which use approximate relations such as Eq. (3). Nevertheless, we see that the upper bound $\omega_{\rm m}^{\rm u}$ is in good agreement with computational results.
@@
In addition to atomic hydrogen with the lightest atomic mass, another appropriate test for the theoretical upper bound $\omega_{\rm m}^{\rm u}$ in Eq. (4) involves the comparison with hydrogen-rich systems. These include hydride superconductors at high pressure, which have been of interest recently because they exhibit the highest superconducting critical temperature $T_{c}$ currently known Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). For our assessment of the maximal frequency upper bound, we chose a range of hydrides based on their chemical and structural variety as well as a wide range of pressures at which phonons have been previously calculated using ab initio methods. Our list includes data on LaH10Durajski et al. (2020) and H3S Sano et al. (2016), with the highest experimentally measured $T_{c}$, two other binary hydrides CaH6Jeon et al. (2022) and SrH10Tanaka et al. (2017), and two ternaries MgIrH6Doliu et al. (2024) and MgVH6Zheng et al. (2021). The maximal phonon frequencies extracted from the calculated dispersion curves of these hydrides are shown in Fig. 2 as red diamonds. We observe that most points are below the fundamental bound $\omega_{\rm m}^{\rm u}$. The maximal frequency in SrH10 at the highest pressure of 300 GPa exceeds the fundamental bound (6) but is below the bound accounting for the pressure effect (7).
@@ 3.1 Upper limit to $T_{c}$
We now discuss the implications of the upper bound of $\omega_{\rm m}$ for phonon-mediated superconductivity and the superconducting critical temperature $T_{c}$. Upper bounds of $T_{c}$ have been discussed on the basis of different mechanisms and have been related to various system properties, including the electron-phonon coupling constant $\lambda$, the energy distribution of valence electrons, the phonon and Fermi energies, phonons and other instabilities, phase fluctuations, and stiffness (see, e.g., Refs. McMillan (1968); Moussa and Cohen (2006); Liu et al. (2022); Varma (2012); Esterlis et al. (2018a, b); Hazra et al. (2019); Chubukov et al. (2020) and references therein). More recent work has offered model examples where these bounds can be exceeded Hofmann et al. (2022).
@@
In the ME picture, $T_{c}=C\,\overline{\omega}$, where $C$ is a slowly-varying monotonically increasing function of $\lambda$ and $\overline{\omega}$ is an averaged phonon frequency Moussa and Cohen (2006); Carbotte (1990). More recent work explored effects beyond the ME theory and noted that they may be inessential for the superconducting state and $T_{c}$ Chubukov et al. (2020). There are several other interesting insights related to the effects of phonons and $\lambda$ on $T_{c}$. First, depending on the system, different phonon modes may contribute differently to $\lambda$ and $T_{c}$. This has been recently discussed in hydrides where different hydrogen-dominated modes combine with other element-dominated modes to set $\lambda$ Durajski et al. (2020); Sano et al. (2016); Tanaka et al. (2017); Jeon et al. (2022); Zheng et al. (2021); Doliu et al. (2024). Second, a more general insight is that $T_{c}$ has a maximum at a certain $C$ due to the competition between the increase of $T_{c}$ with $\lambda$ and the suppression of $T_{c}$ due to instabilities at large $\lambda$ Moussa and Cohen (2006); Esterlis et al. (2018b, a); Carbotte (1990). Supported by model simulations, theory, and empirical data, $C$ is found to be $C\approx 0.1$ Esterlis et al. (2018b, a). A similar $C$ in the range $0.1-0.2$ corresponds to $\lambda=1-2$ Moussa and Cohen (2006); Carbotte (1990), which are typical values of $\lambda$ in hydrides under pressure Shipley et al. (2021) where the highest $T_{c}$ is observed Flores-Livas et al. (2020); Duan et al. (2017); Pickard et al. (2019). We are interested in an upper bound for $T_{c}$, hence we substitute the average $\bar{\omega}$ by the maximal phonon frequency $\omega_{\rm m}$. Then, using Eq. (4) for $\omega_{\rm m}^{\rm u}$, the upper bound of $T_{c}$, which we call $T_{c}^{u}$, is
@@
The range (9) can be verified numerically by calculating an upper bound within the ME theory directly,
@@
where the functional $T_{c}[\alpha^{2}F]$ is the critical temperature from the solution of the Eliasbherg equations Marsiglio (2020) and $S$ is a set of sensible Eliashberg functions extending to a maximum phonon frequency of $\omega_{\rm m}^{\rm u}$. This optimization is shown in Fig. 3 for a fixed value of $\lambda=2$, and converges to $T_{c}^{u}\sim 600$ K, well within the range of Eq. (9). We see that the optimization proceeds by shifting spectral weight to as high a frequency as possible, confirming that the value of $\omega_{m}^{u}$ is the limiting factor. Details of this optimization procedure are available in the Appendix.
@@
$T_{c}^{u}$ in Eq. (8) can be made more stringent in a particular family of materials such as hydrides. Recall that (a) $E_{\rm R}$ used to evaluate the bonding energy $E$ is larger than bonding energies in real materials and (b) the average mass in real materials is larger than the proton mass $m_{p}$ featuring in Eqs. (4) and (8). Both effects reduce $T_{c}^{u}$. According to Eqs. (3), (4), and (8), the difference between $E$ and $E_{\rm R}$ results in $T_{c}^{u}$ decreasing by a factor $\frac{E_{\rm{eff}}}{E_{\rm R}}$, where $E_{\rm{eff}}$ is the effective bonding energy in the relevant subsystem within a hydride. The difference between the average mass and $m_{p}$ reduces $T_{c}^{u}$ by a factor $\left(\frac{m_{\rm{eff}}}{m_{p}}\right)^{\frac{1}{2}}$, where $m_{\rm{eff}}$ is the effective mass in the relevant subsystem. In hydrogen-rich systems, $m_{\rm{eff}}$ is close to $m_{p}$ and this factor is close to one.

### v2

@@ 1 Introduction
The governing role of fundamental physical constants has been extensively discussed in high-energy particle physics and astrophysics. Examples include the stability of nuclei, nuclear reactions, and the formation and evolution of stars – where heavy elements are produced and give rise to other observable structures. In these and other processes, fundamental constants give the Universe its observed properties and separate it from potential other universes where these constants may be different [1, 2, 3, 4, 5, 6, 7, 8, 9]. Understanding the origin of fundamental constants and their fine-tuning is considered one of the grand challenges in science [10, 11]. Approaching this challenge starts with deepening our understanding of how fundamental constants affect phenomena spanning the length and energy scales between particle physics and astrophysics [12, 13, 14].
@@
In condensed matter physics and in other areas that support a continuum description of matter, such as elasticity and hydrodynamics, system properties are driven by many-body collective effects. These effects are often considered as emergent and thus not reducible to individual particles whose properties are set by fundamental constants. It was therefore unexpected when it was discovered that fundamental constants govern the bounds, characteristic values and properties of several key many-body collective processes. Examples include viscosity [15, 16, 17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [13], the speed of sound [28], elastic moduli including those in lower-dimensional systems [13, 12], thermal expansion, melting [29, 30] and velocity gradients that can be set up in cells using biochemical energy [31]. In these examples, properties are governed by either fundamental constants only or by a combination of these constants with other parameters such as temperature. The implications of these findings extend to both fundamental and applied science [12].
@@
Here, we show that fundamental physical constants provide an upper bound to the frequency of phonons in condensed matter phases, setting a limit to the vibrational frequency of atoms. This bound involves the dimensionless electron-to-proton mass ratio $\beta$, the key constant involved in setting the finely-tuned “habitable zone” of our world in the ($\alpha$, $\beta$) space, where $\alpha$ is the fine-structure constant [1]. We find this bound to be in agreement with ab initio simulations of atomic hydrogen and hydride superconductors which, as hydrogen is the lightest nucleus (a single proton), host the highest possible phonon frequencies. We subsequently show that this upper bound in the maximal phonon frequency implies an upper bound to the phonon-mediated superconducting critical temperature $T_{\mathrm{c}}$ in condensed matter. In particular, we show that fundamental constants set the upper bound to $T_{c}$ on the order of $10^{2}-10^{3}$ K, consistent with calculations of $T_{c}$ from optimal Eliashberg functions. We also observe that in addition to electron-phonon coupling, the bound to $T_{c}$ in terms of fundamental constants applies to other coupling mechanisms such as electron-electron coupling. This implies that the current search for $T_{\mathrm{c}}$ at and above $300$ K is itself due to fundamental constants taking their observed values. We finally observe that fundamental constants affect the observability and operation of other phenomena in condensed matter physics including phase transitions. This has wider implications compared to earlier work discussing how fundamental constants set bounds on properties that are already observed [15, 16, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13, 31, 12, 14].
@@ 2 Upper bound on the vibrational frequency
and the Rydberg energy, $E_{\rm R}=\frac{e^{2}}{8\pi\epsilon_{0}a_{\rm B}}$ [32], sets the characteristic scale of $E$ to the order of several eV:
@@
The same ratio (3) follows by combining two known relations in metallic systems: $\frac{\hbar\omega_{\rm m}}{E}\approx\frac{c}{v_{\rm F}}$, where $v_{\rm F}$ is the Fermi velocity and $c$ is the speed of sound, and $\frac{c}{v_{\rm F}}\approx\left(\frac{m_{e}}{m}\right)^{\frac{1}{2}}$, providing an order-of-magnitude estimation of $\frac{\hbar\omega_{\rm m}}{E}$ in other systems too [32].
@@
Equation  (5) offers a simple result showing that the upper bound to the atomic oscillation frequency is set by the Compton frequency times the square of $\alpha$ and the square-root of the electron-to-proton mass ratio.
Equation  (4) sets $\omega_{\rm m}^{\rm u}$ in terms of fundamental physical constants. The constants $\hbar$, $m_{e}$, and $e$ determine the electromagnetic energy scale driving atomic oscillations. The proton mass $m_{p}$ is intrinsic to the atoms being driven and sets the scale (both in time, and space) of the oscillatory response. Interestingly, $\omega_{\rm m}^{\rm u}$ involves the electron-to-proton mass ratio $\beta=\frac{m_{e}}{m_{p}}$. This dimensionless constant is considered of fundamental importance: together with the fine-structure constant, it sets a finely-tuned habitable zone in ($\alpha$, $\beta$) space where it governs the stability of molecular structures, processes involved in igniting and fuelling stars (and thus the production of heavy elements), and other essential effects giving rise to our observable world [1].
@@
Equations  (4)-(6) show that fundamental constants set $\omega_{\rm m}^{\rm u}$ on the order of $10^{3}$$10^{4}$ K, a result we will use to discuss phonon-mediated superconductivity below. The upper bound $\omega_{\rm m}^{\rm u}$ in (6) is of the same order of magnitude, but above, the experimental Debye frequency $\omega_{\rm D}$ of 2240 K in ambient pressure diamond [33], the hardest known material [34] or a similar value of the maximal phonon frequency in graphene [35].
The upper bound (4)-(6) corresponds to solid hydrogen with strong metallic bonding. Although this phase only exists at megabar pressures [36, 37, 38] and is thermodynamically and dynamically unstable at ambient pressure, it is interesting to check the validity of our bound and calculate $\omega_{\rm m}$ in atomic hydrogen. The importance of this calculation is also highlighted by the strong interest in the properties of atomic hydrogen at high pressure (see, e.g., Refs. [36, 37, 38]).
Before calculating $\omega_{\rm m}$, we first evaluate the effect of pressure on the fundamental bound (4) because solid atomic hydrogen exists at high pressure only as mentioned above. We make two remarks in relation to the pressure effect. First, hydrogen is a unique element with no core electrons. As compared to heavier elements, this gives weaker repulsive contributions to the interatomic interaction and weaker pressure dependence of elastic moduli [39]. As a result, pressure has a weaker effect on $\omega_{\rm m}$ in hydrogen as compared to other systems. Second, the effect of pressure on $\omega_{\rm m}^{\rm u}$ can be approximately estimated by adding the term $PV$ to $E$ in Eq. (3), representing the work needed to overcome the elastic deformation [40] due to external pressure in order to surmount the energy of cohesion. Here, $V$ is the elementary volume on the order of $a^{3}$. According to Eqs. (3)-(4), this increases $\omega_{m}^{u}(P)$ as
@@
We note that the full treatment of pressure dependence of frequency for a particular system involves the system-specific equation of state and may involve nonlinear elasticity [41]. This introduces system-specificity in $\omega_{\rm m}^{\rm u}(P)$. Our aim here is to evaluate the effect of pressure in terms of fundamental constants only. This evaluation is necessarily approximate and provides an order-of-magnitude estimation [12, 13]. We will later see that setting $a$ close to $a_{\rm B}$ agrees with the modelling results, showing that our approximate treatment is sensible.
We calculate the phonon dispersion curves of atomic hydrogen in the $I4_{1}/amd$ structure [42, 43], which is currently the best candidate structure for solid atomic metallic hydrogen. This structure is calculated to become thermodynamically stable in the pressure range 400400400400600600600600 GPa [44, 45, 46], below which solid hydrogen is a molecular solid. We perform density functional theory (DFT) calculations using the castep package [47]. Details of DFT calculations are given in the Appendix.
@@
Figure 2: Maximal phonon frequency as a function of pressure for different systems. Blue bullet points show the calculated maximal phonon frequencies in atomic hydrogen at the corresponding pressures. Red diamonds show the maximal phonon frequencies in a range of hydrides at different pressures, which are, from left to right: MgIrH6 [48], MgVH6 [49], CaH6 [50], LaH10 (top) [51], H3S (bottom) [52], and SrH10 [53]. The horizontal solid black line shows the fundamental bound $\omega_{\rm m}^{\rm u}$ (6). The horizontal dash-dotted line shows the Debye frequency of diamond at ambient pressure. The dashed line on top of the graph shows $\omega_{\rm m}^{\rm u}(P)$ in Eq. (7) with $a=2a_{\rm B}$.
@@
We observe in Fig. 2 that the calculated $\omega_{\rm m}$ is only slightly above $\omega_{\rm m}^{\rm u}$ at the high pressure of 400 GPa in atomic hydrogen. Assuming a pressure-independent effective $a$ in Eq. (7) for estimation purposes, we find that $\omega_{\rm m}^{\rm u}(P)$ lies close and above the hydrogen points if $a$ is taken to be about $2a_{\rm B}$ (see dashed black line in Fig. 2). This is consistent with the elementary volume being related to the Bohr radius and fundamental constants. Although $a=2a_{\rm B}$ in Eq. (7) introduces an ad hoc choice, the dashed line in Fig. 2 demonstrates that the pressure effect on $\omega_{\rm m}^{\rm u}$ can be accounted for with a sensible and physically justified choice of $a$. In this regard, we note more generally that evaluating bounds on many important condensed matter properties using fundamental constants only inevitably involves approximations [12, 13]. This includes $\omega_{\rm m}^{\rm u}$ in Eqs. (4)-(6) which use approximate relations such as Eq. (3). Nevertheless, we see that the upper bound $\omega_{\rm m}^{\rm u}$ is in good agreement with computational results.
@@
In addition to atomic hydrogen with the lightest atomic mass, another appropriate test for the theoretical upper bound $\omega_{\rm m}^{\rm u}$ in Eq. (4) involves the comparison with hydrogen-rich systems. These include hydride superconductors at high pressure, which have been of interest recently. These systems exhibit the highest superconducting critical temperature $T_{c}$ currently known extending to 250 K [54, 55, 56]. For our assessment of the maximal frequency upper bound, we chose a range of hydrides based on their chemical and structural variety as well as a wide range of pressures at which phonons have been previously calculated using ab initio methods. Our list includes data on LaH10[51] and H3S [52], with the highest experimentally measured $T_{c}$, two other binary hydrides CaH6[50] and SrH10[53], and two ternaries MgIrH6[48] and MgVH6[49]. The maximal phonon frequencies extracted from the calculated dispersion curves of these hydrides are shown in Fig. 2 as red diamonds. We observe that most points are below the fundamental bound $\omega_{\rm m}^{\rm u}$. The maximal frequency in SrH10 at the highest pressure of 300 GPa exceeds the fundamental bound (6) but is below the bound accounting for the pressure effect (7).
@@ 3.1 Upper limit to $T_{c}$
We now discuss the implications of the upper bound of $\omega_{\rm m}$ for phonon-mediated superconductivity and the superconducting critical temperature $T_{c}$. Upper bounds of $T_{c}$ have been discussed on the basis of different mechanisms and have been related to various system properties, including the electron-phonon coupling constant $\lambda$, the energy distribution of valence electrons, the phonon and Fermi energies, instabilities involving phonons, electrons and other instability types, phase fluctuations, and stiffness (see, e.g., Refs. [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] and references therein). More recent work has offered model examples where these bounds can be exceeded [67].
@@
In the ME picture, $T_{c}=C\,\overline{\omega}$, where $C$ is a monotonically increasing function of $\lambda$ and $\overline{\omega}$ is an averaged phonon frequency [58, 68]. More recent work explored effects beyond the ME theory and noted that they may be inessential for the superconducting state and $T_{c}$ [64]. There are several other interesting insights related to the effects of phonons and $\lambda$ on $T_{c}$. First, depending on the system, different phonon modes may contribute differently to $\lambda$ and $T_{c}$. This has been recently discussed in hydrides where different hydrogen-dominated modes combine with other element-dominated modes to set $\lambda$ [51, 52, 53, 50, 49, 48]. Second, a more general insight is that $T_{c}$ has a maximum at a certain $C$ due to the competition between the increase of $T_{c}$ with $\lambda$ and the suppression of $T_{c}$ due to instabilities at large $\lambda$ [58, 62, 61, 68, 69] (see Ref. [70] of the effect of soft phonons in this regard). Supported by model simulations, theory, and empirical data, $C$ is $C\approx 0.1$ [62, 61] and increases to 0.2-0.3 for a different type of instability [66]. Assuming that different types of instabilities all operate, the upper bound to $T_{c}$ corresponds to the lower range of $C$ which induces the instability and re-organisation first, with the associated lowering of $T_{c}$. Interestingly, a similar $C$ in the range $0.1-0.2$ corresponds to $\lambda=1-2$ [58, 68], which are typical values of $\lambda$ in hydrides under pressure [71] where the highest $T_{c}$ is observed [54, 55, 56]. We are interested in an upper bound for $T_{c}$, hence we substitute the average $\bar{\omega}$ by the maximal phonon frequency $\omega_{\rm m}$. Then, using Eq. (4) for $\omega_{\rm m}^{\rm u}$ at zero pressure, the upper bound of $T_{c}$, which we call $T_{c}^{u}$, is
@@
The range (9) can be verified numerically from ME theory, where the central quantity is the phonon-frequency-resolved coupling strength, given by the Eliashberg function $\alpha^{2}F(\omega)$. Restricting $\alpha^{2}F(\omega)$ to frequencies below the upper bound $\omega_{m}^{u}$ (6), we can directly calculate an upper bound on $T_{c}$ as
@@
where the functional $T_{c}[\alpha^{2}F]$ is the critical temperature from the solution of the Eliashberg equations [72] and $S$ is a set of sensible Eliashberg functions obeying the upper bound on $\omega$, with a $\lambda$ value limited to avoid instabilities as discussed earlier [58, 62, 61, 68, 69]. Details of this optimization procedure are available in the Appendix.
Such an optimization is shown in Fig. 3 for a fixed value of $\lambda=2$, and converges to $T_{c}^{u}\sim 600$ K, well within the range of Eq. (9). Since $T_{c}$ is the result of a trade-off between $\lambda$ and $\bar{\omega}$ [70] one would expect that, in the face of a $\lambda$ limited by instability, $\bar{\omega}$ becomes the central quantity. Indeed, we see that the optimization proceeds by shifting spectral weight to as high a frequency as possible, corroborating other work employing dispersionless (Einstein) phonons to obtain an upper bound to $T_{c}$ [66], and confirming that the value of $\omega_{m}^{u}$ is the limiting factor in the optimization.
@@
$T_{c}^{u}$ in Eq. (8) can be made more stringent in a particular family of materials. Recall that (a) $E_{\rm R}$ used to evaluate the bonding energy $E$ is larger than bonding energies in real materials and (b) the average mass in real materials is larger than the proton mass $m_{p}$ featuring in Eqs. (4) and (8). Both effects reduce $T_{c}^{u}$. According to Eqs. (3), (4), and (8), the difference between $E$ and $E_{\rm R}$ results in $T_{c}^{u}$ decreasing by a factor $\frac{E_{\rm{eff}}}{E_{\rm R}}$, where $E_{\rm{eff}}$ is the effective bonding energy in the relevant subsystem. The difference between the average mass and $m_{p}$ reduces $T_{c}^{u}$ by a factor $\left(\frac{m_{\rm{eff}}}{m_{p}}\right)^{\frac{1}{2}}$, where $m_{\rm{eff}}$ is the effective mass in the relevant subsystem. The same reasoning applies to the upper frequency bound (4).
We note that we do not consider the effect of pressure on $T_{c}$. $T_{c}$ can be a non-monotonic function of pressure and involve complex effects including, for example, effects of pressure on phonon frequencies, $\lambda$, screened Coulomb pseudopotential and so on [73]. In particular, to repeat the analysis in this work, the affect of pressure on the instability limit to $\lambda$ would need to be accounted for.
We also note that our bound (8) invited a recent discussion in Ref. [74]. This work derived an upper bound to $T_{c}$ using different arguments and interestingly showed that this bound is expressed as the same combination of fundamental constants as Eq. (8).